OpenStudy (anonymous):

Solve: x^4-13x^2+36=0 A. -3,-2,2,3 B.-9,-4,4,9 C.2,3,2i,3i D.-2,-3,2i,3i

OpenStudy (anonymous):

last question tonight I swear! =D

OpenStudy (lgbasallote):

it seems like a case of quadratics...let a = x^2 so a^2 = x^4 a^2 - 13a + 36 does that make it easier?

OpenStudy (anonymous):

Oh yeah, I remember now! Thank you so much!

OpenStudy (lgbasallote):

dont forget to sub back to x^2 when you're done

OpenStudy (anonymous):

first you have to factor x^4-13x^2+36 into two brackets... so first off, you want to draw your two brackets...( )( ) ...then it is just trial and error to get the right order of the terms inside the brackets to get you the original equation... now, your equation starts off with an x^4, not x^2, so you have to be careful of what you put in each bracket first